For an economy without government and foreign trade with MPS=0.20 Describe the multiplier process if autonomous investment increases by 10 units.
1
Expert's answer
2020-06-24T13:43:24-0400

We compute the multiplier as:


"m = \dfrac{\Delta Y}{\Delta I} = \dfrac{1}{MPS}"

If MPS = 0.20, then:



"\dfrac{\Delta Y}{\Delta I} = \dfrac{1}{0.2} = 5"

A multiplier of 5 means that the equilibrium income would increase by 5 units when the autonomous investment increases by 1 unit.